A mortar shell is s feet above the ground after t seconds, where s = f(t) = -16t^2 + 512t + 64. Find the height of the shell 30 seconds after it is fired. (Simplify the answer as much as possible.)

Could someone help me with this problem?
Thank you.

Answer by richwmiller(17219)

plug in 30 for t
s=-16t^2 + 512t + 64
s = -16 (t^2-32 t-4)
s = -16*(30^2-32(30)-4)
s = -16(-2(30)-4)
s=-16*(-64)
s=1024